evalute f (-3) if f(x)=-x^2+2
f(2) if f(x)=-5x+6
A function f is represented verbally by "cube the input x and then add 4".
give the symbolic representation of f.
Determin if f (x)=-5/2+3x is a linear function, yes it is.. give value for a and b so that f may be expressed as f(x)=ax+b

Answer by stanbon(75887)

evalute f (-3) if f(x)=-x^2+2
f(-3) = -(-3)^2 + 2 = -9+2 = -7
------------------------------------
f(2) if f(x)=-5x+6
f(2) = -5*2+6 = -10+6 = -4
------------------------------------
A function f is represented verbally by "cube the input x and then add 4".
give the symbolic representation of f.
f(x) = x^3+4
------------------------------------
Determin if f (x)=-5/2+3x is a linear function, yes it is.. give value for a and b so that f may be expressed as f(x)=ax+b
---
f(x) = 3x -(5/2)
a = 3 ; b = -5/2
==============================
Cheers,
Stan H.
=======================================